# publications.bib

### v3

@@ 3. Mantovan’s Formula and the Hodge-Tate Period Morphism
$\mathcal{S}_{K^{p},C}^{b,\circ,\mathrm{rd}}\times_{\mathcal{F}\ell_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b}\hookrightarrow\mathcal{S}^{b,\circ}_{K^{p},C}$
is a qcqs open immersion, which is an isomorphism on rank $1$ points. In turn, it induces an isomorphism of canonical compactifications over the partially proper strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b}$. Therefore, by passing to canonical compactifications over $\mathcal{F}\ell_{G,\mu^{-1}}^{b}$, we deduce that diagram (10) is also Cartesian. ∎
We now invoke a result of Zhang [Zha].
@@
[Zha, Theorem 1.3] Assuming 1.10, for all $b\in B(G,\mu)$ the Cartesian diagram (10) extends to a Cartesian diagram
(13) ${\left[\mathcal{S}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}\right]}$${\left[\mathcal{F}\ell^{b}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]}$${\left[\overline{\mathfrak{Ig}}_{C}^{b,*}/\mathcal{J}_{b}\right]}$${\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right]}$$\scriptstyle{\pi_{\mathrm{HT}}^{b,*}}$$\scriptstyle{\phantom{}{}^{*}\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{h^{\leftarrow}_{b}}$$\scriptstyle{\pi_{\mathfrak{Ig}}^{b,*}}$
@@
The rough idea behind proving this is to apply a relative $\mathop{\rm Spa}$ construction in the category of diamonds to the horizontal maps of the diagram (13) by invoking Hartogs’ principle, as in Proposition 3.8.
We now state the key Corollary that we will need.
@@
Assuming 1.10, for all $b\in B(G,\mu)$ we have a Cartesian diagram
(14) ${\left[\mathcal{S}^{b}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}\right]}$${\mathcal{F}\ell^{b}_{G,\mu^{-1}}}$${\left[(\overline{\mathfrak{Ig}}_{C}^{b,*}\setminus\overline{\partial\mathfrak{Ig}}_{C}^{b,*})/\mathcal{J}_{b}\right]}$${\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right].}$$\scriptstyle{\pi_{\mathrm{HT}}^{b}}$$\scriptstyle{\phantom{}{}^{\partial}\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{h^{\leftarrow}_{b}}$$\scriptstyle{\pi_{\mathfrak{Ig}}^{b,\partial}}$
@@
This follows from the Cartesian diagram (13) and Corollary 3.5. ∎
@@
$i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$
we can see that this is canonically identified with
$h_{b}^{\leftarrow*}R\pi_{\mathfrak{Ig}!}^{b,\partial}(\overline{\mathbb{F}}_{\ell})$
via proper base change. Moreover, we can identify $R\pi_{\mathfrak{Ig}!}^{b,\partial}(\overline{\mathbb{F}}_{\ell})$ simply with the complex $V_{b}:=R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ of $J_{b}(\mathbb{Q}_{p})$-modules under the identification $\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, as in Corollary 3.6. We can further refine this using the following lemma.
@@
$i_{b*}h_{b}^{\leftarrow*}(V_{b})\simeq h^{\leftarrow*}j_{b*}(V_{b})$
of sheaves on $\mathcal{F}\ell_{G,\mu^{-1}}$.
@@
The first isomorphism follows from proper base-change. For the second isomorphism, we note that
$h^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow\mathrm{Bun}_{G}$
is cohomologically smooth separated and representable in locally spatial diamonds; therefore, the result follows by smooth base-change [Ecod, Proposition 23.16 (2)]. ∎
@@
$R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}],i_{b!}i_{b}^{*}(R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})))$
on the cohomology of the Shimura variety are identified with $h^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})\in\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, and similarly for $R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}],i_{b*}i_{b}^{*}(R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})))$ and $h^{\rightarrow}_{*}h^{\leftarrow*}j_{b*}(V_{b})$.
@@
Assuming 1.10, we have a filtration on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ by complexes of smooth representations of $G(\mathbb{Q}_{p})$, with graded pieces isomorphic to $h^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})$, where $V_{b}\simeq R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$.
@@
[…]
where $\mathrm{Hck}_{G,\leq\mu}$ is the stack parametrizing modifications $\mathcal{E}_{1}\rightarrow\mathcal{E}_{2}$ of a pair of $G$-bundles with meromorphy bounded by $\mu$ over the fixed untilt defined by $C$. We define the Hecke operator [FS, Section IX.2]
@@
where $E_{\mu}$ is the reflex field of $\mu$ and $\mathcal{S}_{\mu}$ is a sheaf on $\mathrm{Hck}_{G,\leq\mu}$ attached to the highest weight tilting module $\mathcal{T}_{\mu}$ by geometric Satake [^4] We note that, using [FS, Proposition VII.5.2], we can replace the natural push-forward in the category of solid sheaves with the * push-forward in the usual category of étale $\overline{\mathbb{F}}_{\ell}$-sheaves when defining the Hecke operator.. Here $E_{\mu}$ denotes the reflex field of $\mu$.
If we now let $\mu$ be the minuscule cocharacter appearing above then the Bialynicki-Birula map gives an isomorphism of diamonds between the open locus of $\mathrm{Hck}_{G,\leq\mu}$ where $\mathcal{E}_{1}$ is isomorphic to the trivial bundle and $[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$, which identifies $h_{\mu}^{\rightarrow}$ (resp. $h_{\mu}^{\leftarrow})$) with $h^{\rightarrow}$ (resp. $h^{\leftarrow}$). Moreover, this is a cohomologically smooth space of dimension $d:=\langle 2\rho_{G},\mu\rangle$, and we have an isomorphism $\mathcal{S}_{\mu}\simeq\overline{\mathbb{F}}_{\ell}[d](\frac{d}{2})$ [^5] This is true for the highest weight module $V_{\mu}$ and this agrees with the highest weight tilting module $\mathcal{T}_{\mu}$, since $\mu$ is minuscule.. It follows, by proper base-change, that we have an isomorphism
$h^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})\simeq j_{1}^{*}T_{\mu}(j_{b!}(V_{b}))[-d](-\frac{d}{2})$
of $G(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-modules, where $1\in B(G)$ is the trivial element. The $W_{E_{\mathfrak{p}}}$-equivariance follows since the above Cartesian diagrams all descend to $\breve{E}_{\mathfrak{p}}$ and are also compatible with the Frobenius descent datum on $\mathrm{Sht}(G,b,\mu)_{\infty,C}\rightarrow\mathop{\rm Spd}(\breve{E}_{\mathfrak{p}})$ (This is true for (9) by the results of [CS1] and all the other diagrams are constructed from this one). This gives us Theorem 1.12.
@@
Assuming 1.10, the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ has a $G(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-equivariant filtration with graded pieces given by $j_{1}^{*}T_{\mu}(j_{b!}(V_{b}))[-d](-\frac{d}{2})$ for varying $b\in B(G,\mu)$, where $V_{b}\simeq R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$.
@@
as $G(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-modules. Here
@@
is a complex of $G(\mathbb{Q}_{p})\times J_{b}(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-modules, where $\mathrm{Sht}(G,b,\mu)_{\infty,C}$ is as defined above, and $\overline{\mathbb{F}}_{\ell}(d_{b})$ is the sheaf with trivial Weil group action and $J_{b}(\mathbb{Q}_{p})$-action given as in [Ko, Lemma 7.4].
@@
It remains to explain the description of $j_{1}^{*}T_{\mu}(j_{b!}(V_{b}))[-d](-\frac{d}{2})$. By applying the second part of [Ham2, Proposition 11.12] and noting that $\mathcal{S}_{\mu}\simeq\overline{\mathbb{F}}_{\ell}[d](\frac{d}{2})$ since $\mu$ is minuscule (where we recall that, since $\mu$ is minuscule, the representation $\mathcal{T}_{\mu}$ agrees with the usual highest weight representation), we obtain that the graded pieces are isomorphic to
@@
as desired, where $\kappa$ is the character of $J_{b}(\mathbb{Q}_{p})$ defined by the action of $J_{b}(\mathbb{Q}_{p})$ on the compactly supported cohomology of the $\ell$-adically contractible group diamond $\mathcal{J}_{b}^{>0}$, where $\mathcal{J}_{b}\simeq\mathcal{J}_{b}^{>0}\ltimes J_{b}(\mathbb{Q}_{p})$ is the semi-direct product structure given by allowing $\mathrm{Aut}(\mathcal{E}_{b})$ to act on its canonical reduction. However, by combining this with [Ko, Lemma 7.6] and its proof, we can rewrite this as

### v5

@@ 3.2. Mantovan’s Product Formula
$\mathcal{S}_{K^{p},C}^{b,\circ,\mathrm{rd}}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}\hookrightarrow\mathcal{S}^{b,\circ}_{K^{p},C}$
is a qcqs open immersion, which is an isomorphism on rank $1$ points. In turn, it induces an isomorphism of canonical compactifications over the partially proper strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$. Therefore, by passing to canonical compactifications over $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$, we deduce that the diagram (14) is also Cartesian, using that the formation of canonical compactifications commutes with limits [56, Proposition 18.7 (v)]. ∎
Remark 3.13.
We note that our conventions for moduli spaces of shtukas differs from that in [55, Definition 23.1.1]; in particular, there the space $\mathrm{Sht}(G,b,\mu)_{\infty}$ denotes the moduli space $\mathcal{E}_{0}\dashrightarrow\mathcal{E}_{b}$ parametrizes modifications of meromorphy $\mu$, which would identify more naturally with the space we denote by $\mathrm{Sht}(G,b,\mu^{-1})_{\infty}$. This is consistent with [27, 24, 47], where this convention is nice in that the Hecke operator $j_{1}^{*}T_{\mu}$ is related to $\mathrm{Sht}(G,b,\mu)_{\infty}$ for $b\in B(G,\mu)$ instead of $\mathrm{Sht}(G,b,\mu^{-1})$, as in the usual convention. Here we also note that we have uniformized the Schubert stratification of the $\mathbb{B}_{\mathrm{dR}}^{+}$-Grassmannian so that $G(\mathbb{B}_{\mathrm{dR},C}):=\bigsqcup_{\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}}G(\mathbb{B}^{+}_{\mathrm{dR},C})\mu(\xi)^{-1}G(\mathbb{B}_{\mathrm{dR},C}^{+})$, where the coset $G(\mathbb{B}^{+}_{\mathrm{dR},C})\mu(\xi)^{-1}G(\mathbb{B}_{\mathrm{dR},C}^{+})$ defines $\mathrm{Gr}_{G,\mu}$, as in the discussion proceeding [7, Definition 3.4.1]. We warn the reader that this differs by an inverse from the uniformization used in [55] and [10], which is why the index set for the adic Newton stratification of $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ is for us $B(G,\mu)$ instead of $B(G,\mu^{-1})$.
We now invoke a result of Zhang [68].
@@
[68, Theorem 1.3, Section 9.4] Assuming 1.13, for all $b\in B(G,\mu)$ the Cartesian diagram (14) extends to a Cartesian diagram
(17) $\left[\mathcal{S}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\mathcal{F}\ell^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\overline{\mathfrak{Ig}}_{C}^{b,*}/\mathcal{J}_{b}\right]$$\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right]$$\scriptstyle{\pi_{\mathrm{HT}}^{b,*}}$$\scriptstyle{\phantom{}{}^{*}\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{h^{\leftarrow}_{b}}$$\scriptstyle{\pi_{\mathfrak{Ig}}^{b,*}}$
@@
The rough idea behind proving this is to apply a relative $\mathop{\rm Spa}$ construction in the category of $v$-sheaves to the horizontal maps of the diagram (17) by invoking Hartogs’ principle, as in Proposition 3.9.
We now state the key Corollary that we will need. We write $\tilde{s}_{b}:[\widetilde{\mathcal{F}\ell}_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ for the $\mathcal{J}_{b}^{>0}$-torsor defined by pulling back $h^{\leftarrow}_{b}$ along the natural map $s_{b}:[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$ induced by (10). We write $\tilde{\pi}_{\mathrm{HT}}^{b}:[\widetilde{\mathcal{S}}^{b}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\widetilde{\mathcal{F}\ell}_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ and $\tilde{\pi}_{\mathrm{HT}}^{b,*}:[\widetilde{\mathcal{S}}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\widetilde{\mathcal{F}\ell}^{\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$ for the pullbacks of $\pi_{\mathrm{HT}}^{b}$ and $\pi_{\mathrm{HT}}^{b,*}$ along the map $\widetilde{\mathcal{F}\ell}_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$.
We now have the following.
@@
Assuming 1.13, for all $b\in B(G,\mu)$ we have a Cartesian diagram
(18) $\left[\widetilde{\mathcal{S}}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\widetilde{\mathcal{F}\ell}^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[(\overline{\mathfrak{Ig}}_{C}^{b,*}\setminus\overline{\partial\mathfrak{Ig}}_{C}^{b,*})/\underline{J_{b}(\mathbb{Q}_{p})}\right]$$\left[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}\right].$$\scriptstyle{\tilde{\pi}_{\mathrm{HT}}^{b,*}}$$\scriptstyle{\phantom{}{}^{\partial}\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{\tilde{\pi}_{\mathfrak{Ig}}^{b,\partial}}$
with maps as defined above
@@
This follows from the Cartesian diagram (17) and Corollary 3.4. More precisely, consider the open substack
$\widetilde{\mathcal{S}}_{K^{p},C}^{b,*}/\underline{G(\mathbb{Q}_{p})}]\times_{[\overline{\mathfrak{Ig}}^{b,*}_{C}/\underline{J_{b}(\mathbb{Q}_{p})}]}\left[(\overline{\mathfrak{Ig}}_{C}^{b,*}\setminus\overline{\partial\mathfrak{Ig}}_{C}^{b,*})/\underline{J_{b}(\mathbb{Q}_{p})}\right]\hookrightarrow[\widetilde{\mathcal{S}}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]$
obtained by considering the base-change of $\left[\widetilde{\mathcal{F}\ell}^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]\rightarrow[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]$ along the open immersion $\overline{\mathfrak{Ig}}_{C}^{b,*}\setminus\overline{\partial\mathfrak{Ig}}^{b,*}_{C}/\underline{J_{b}(\mathbb{Q}_{p})}]\hookrightarrow[\overline{\mathfrak{Ig}}^{b,*}_{C}/\underline{J_{b}(\mathbb{Q}_{p})}]$. The claim follows from showing that this open substack identifies with the open substack
$\widetilde{\mathcal{S}}^{b}_{K_{p},C}/\underline{G(\mathbb{Q}_{p})}]\hookrightarrow[\widetilde{\mathcal{S}}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}].$
Using [56, Proposition 11.15], it suffices to check that the open subspaces of the underlying topological space of $\widetilde{\mathcal{S}}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]$ that these two substacks define agree with one another. Since both subspaces are partially proper (as they are $\mathcal{J}_{b}^{>0}$-torsors over something partially proper), it suffices to show that they have the same rank $1$ points, and therefore it suffices to show that they identify with the same subspace after taking the fiber of $\tilde{\pi}_{\mathrm{HT}}^{b,*}$ over a rank $1$ point of $\widetilde{\mathcal{F}\ell}_{G,\mu^{-1}}^{b,\Diamond}$, and this reduces us to Corollary 3.4. ∎
@@
$\tilde{s}_{b}^{*}i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$
we can see that this is canonically identified with $R\tilde{\pi}_{\mathfrak{Ig}!}^{b,\partial}(\overline{\mathbb{F}}_{\ell})\in\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ pulled back along the natural map $\widetilde{\mathcal{F}\ell}_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]$. Moreover, we can identify $R\tilde{\pi}_{\mathfrak{Ig}!}^{b,\partial}(\overline{\mathbb{F}}_{\ell})$ simply with the complex $V_{b}:=R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ of $J_{b}(\mathbb{Q}_{p})$-modules, as in Corollary 3.5. We recall that we have an identification $\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, which is induced via pullback along $p_{b}:\mathrm{Bun}_{G}^{b}\simeq[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\rightarrow[\ast/\underline{J_{b}(\mathbb{Q}_{p})}]$, as in [16, Proposition V.2.2]. The map $s_{b}$ is a section to $p_{b}$ induced by the splitting of the Harder-Narasimhan filtration and $s_{b}^{*}$ is inverse to $p_{b}^{*}$. If we write $\tilde{p}_{b}:[\mathcal{F}\ell^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\widetilde{\mathcal{F}\ell}^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$ for the map obtained by the base-change of $p_{b}$ then $\tilde{s}_{b}$ will be a section of $\tilde{p}_{b}$. In particular, we conclude that
$\tilde{p}_{b}^{*}\tilde{s}_{b}^{*}i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})\simeq i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$
is the pullback of $V_{b}$ along the map $\widetilde{\mathcal{F}\ell}^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]$. Abusively identifying $V_{b}$ and $p_{b}^{*}(V_{b})$ as in the identification $\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, this allows us to deduce an isomorphism
$h^{\leftarrow*}_{b}(V_{b})\simeq i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell}).$
We can further refine this using the following Lemma.
@@
$i_{b*}h_{b}^{\leftarrow*}(V_{b})\simeq h^{\leftarrow*}Rj_{b*}(V_{b})$
of sheaves on $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$.
@@
The first isomorphism follows from proper base-change ([56, Theorem 1.9 (ii)]). For the second isomorphism, we note that
$h^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow\mathrm{Bun}_{G,C}$
is cohomologically smooth separated and representable in locally spatial diamonds; therefore, the result follows by smooth base-change [56, Proposition 23.16 (2)]. ∎
@@
$R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}],i_{b!}i_{b}^{*}(R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})))$
on the cohomology of the Shimura variety coming from excision applied to $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ and the Newton stratification of the flag variety are identified with $Rh^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})\in\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, and similarly for $R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}],i_{b*}i_{b}^{*}(R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})))$ and $Rh^{\rightarrow}_{*}h^{\leftarrow*}Rj_{b*}(V_{b})$.
@@
Assuming 1.13, we have a filtration on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ by complexes of smooth representations of $G(\mathbb{Q}_{p})$, with graded pieces isomorphic to $Rh^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})$, where $V_{b}:=R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})\in\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ is regarded as a sheaf on $\mathrm{Bun}_{G}^{b}$ via the natural identificaiton $\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$.
@@
$\mathrm{Hck}_{G,\leq\mu}$$\mathrm{Bun}_{G}$$\mathrm{Bun}_{G}\times\mathop{\rm Spd}(C)$$\scriptstyle{h_{\mu}^{\leftarrow}}$$\scriptstyle{h_{\mu}^{\rightarrow}}$
where $\mathrm{Hck}_{G,\leq\mu}$ is the stack parametrizing modifications $\mathcal{E}_{1}\rightarrow\mathcal{E}_{2}$ of a pair of $G$-bundles with meromorphy bounded by $\mu$ over the fixed untilt defined by $C$. We define the Hecke operator [16, Section IX.2]
@@
$A\mapsto h_{\mu*}^{\rightarrow}(h_{\mu}^{\leftarrow*}(A)\otimes^{\mathbb{L}}\mathcal{S}_{\mu}),$
where $E_{\mu}$ is the reflex field of $\mu$ and $\mathcal{S}_{\mu}$ is a sheaf on $\mathrm{Hck}_{G,\leq\mu}$ attached to the highest weight tilting module $\mathcal{T}_{\mu}$ by geometric Satake.
Remark 3.21.
To see why this operator gives a functor of the desired shape, we note that we have an isomorphism $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(\mathrm{Bun}_{G}\times\mathop{\rm Spd}(C),\overline{\mathbb{F}}_{\ell})$ by [16, Corollary V.2.3.], and, under this isomorphism, the operator $T_{\mu}$ a priori only defines a functor $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$; however, it follows from [16, Corollary IX.2.3] and the surrounding discussion that this factors through the forgetful functor $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ forgetting the continuous $W_{E_{\mu}}$-action.
If we now let $\mu$ be the minuscule cocharacter appearing above then the Bialynicki-Birula map gives an isomorphism of diamonds between the open locus of $\mathrm{Hck}_{G,\leq\mu}$ where $\mathcal{E}_{1}$ is isomorphic to the trivial bundle and $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$, which identifies $h_{\mu}^{\rightarrow}$ (resp. $h_{\mu}^{\leftarrow})$) with $h^{\rightarrow}$ (resp. $h^{\leftarrow}$). Moreover, this is a cohomologically smooth space of dimension $d:=\langle 2\rho_{G},\mu\rangle$, and we have an isomorphism $\mathcal{S}_{\mu}\simeq\overline{\mathbb{F}}_{\ell}[d](\frac{d}{2})$ [^8] This is true for Satake sheaf attached to the highest weight module $V_{\mu}$ by definition and this agrees with the highest weight tilting module $\mathcal{T}_{\mu}$, since $\mu$ is minuscule.. It follows, by proper base-change, that we have an isomorphism
$h^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})\simeq j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})$
of $G(\mathbb{Q}_{p})$-modules, where $1\in B(G)$ denotes the trivial element. This gives us Theorem 1.15.
@@
Assuming 1.13, the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ has a $G(\mathbb{Q}_{p})$-equivariant filtration with graded pieces given by $j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})$ for varying $b\in B(G,\mu)$, where $V_{b}\simeq R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ and we implicitly identify $T_{\mu}$ with its postcomposition with the forgetful functor $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{p})^{BW_{E_{\mathfrak{p}}}}\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{p})$.
@@
as $G(\mathbb{Q}_{p})$-modules. Here
@@
is a complex of $G(\mathbb{Q}_{p})\times J_{b}(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-modules, where $\mathrm{Sht}(G,b,\mu)_{\infty,C}$ is as defined above, and $\overline{\mathbb{F}}_{\ell}(d_{b})$ is the sheaf with $J_{b}(\mathbb{Q}_{p})$-action given as in [37, Lemma 7.4].
@@
It remains to explain the description of $j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})$. By applying the second part of [24, Proposition 10.3] and noting that $\mathcal{S}_{\mu}\simeq\overline{\mathbb{F}}_{\ell}[d](\frac{d}{2})$ since $\mu$ is minuscule (where we recall that, since $\mu$ is minuscule, the representation $\mathcal{T}_{\mu}$ agrees with the usual highest weight representation), we obtain that the graded pieces are isomorphic to
@@
as desired, where $\kappa$ is the character of $J_{b}(\mathbb{Q}_{p})$ defined by the action of $J_{b}(\mathbb{Q}_{p})$ on the compactly supported cohomology of the $\ell$-adically contractible group diamond $\mathcal{J}_{b}^{>0}$, where $\mathcal{J}_{b}\simeq\mathcal{J}_{b}^{>0}\ltimes J_{b}(\mathbb{Q}_{p})$ is the semi-direct product structure given by allowing $\mathrm{Aut}(\mathcal{E}_{b})$ to act on its canonical reduction. However, by combining this with [37, Lemma 7.6] and its proof, we can rewrite this as